# Dial-In LLM: Human-Aligned LLM-in-the-loop Intent Clustering for Customer Service Dialogues

### v2

@@ Abstract
Discovering customer intentions in dialogue conversations is crucial for automated service agents. Yet, existing intent clustering methods often fail to align with human perceptions due to the heavy reliance on embedding distance metrics and sentence embeddings. To address these limitations, we propose integrating the semantic understanding capabilities of LLMs into an LLM-in-the-loop (LLM-ITL) intent clustering framework. Specifically, this paper (1) investigates the effectiveness of fine-tuned LLMs in semantic coherence evaluation and intent cluster naming, achieving over 95% accuracy; (2) designs an LLM-ITL clustering algorithm that facilitates the iterative discovery of coherent intent clusters; and (3) proposes task-specific techniques tailored for customer service dialogue intent clustering. Since existing English benchmarks pose limited semantic diversity and intent labels, we introduced a comprehensive Chinese dialogue intent dataset, comprising over 100,000 real customer service calls and 1,507 human-annotated intent clusters. The proposed approaches significantly outperformed LLM-guided baselines, achieving notable improvements in clustering quality and a 12% boost in the downstream intent classification task. Combined with several best practices, our findings highlight the potential of LLM-in-the-loop techniques for scalable and human-aligned problem-solving. Sample code and datasets are available at: anonymous.4open.science/r/Dial-in-LLM.
@@
Mengze Hong1  Di Jiang2thanks: Corresponding author: dijiang@webank.com  Yuanfeng Song2  Lu Wang3  Wailing Ng1 Yanjie Sun1  Chen Jason Zhang1  Qing Li1 1Hong Kong Polytechnic University  2AI Group, WeBank Co., Ltd  3Shenzhen University
@@ 1 Introduction
Intent discovery is an essential task in many NLP applications, such as designing AI-driven dialogue systems Hengst et al. (2024) and learning utterance patterns Ghosal et al. (2020). These systems often rely on an intent classification model that assigns intent labels. However, obtaining labeled data for training such supervised models is difficult, especially for dialogue data where topics evolve and contexts change Hengst et al. (2024); Hong et al. (2024). This motivates the need for robust intent discovery methods that can adapt to dynamic dialogue variations and applications.
Figure 1: Comparison of embedding-distance metric and human perception: cosine similarity failed to identify the same intentions under linguistic diversity (top), as well as distinguish different intentions under semantic diversity (bottom).
Intent clustering, which has evolved from unsupervised text clustering, offers significant advantages in automatically discovering thematic relationships Gung et al. (2023). Although text clustering is widely studied, existing research mainly concentrates on developing meaningful representations for text sentences Yin et al. (2021), treating it as a conventional clustering problem without addressing distinctive characteristics such as linguistic and semantic diversity (see Figure 1). Moreover, clustering algorithms heavily rely on quantitative optimization of embedding distance metrics with encoded representations, which is believed to hinder human-interpretable evaluation and validation.
With the growing research interest in integrating Large Language Models (LLMs) into the problem-solving pipeline Zhang et al. (2023); Viswanathan et al. (2024), known as the LLM-in-the-loop (LLM-ITL) paradigm Hong et al. (2025), this paper proposes an LLM-ITL solution specifically designed for dialogue intent clustering. This approach emphasizes the use of intermediate clustering results with human-aligned LLM utilities, providing practical benefits to industrial applications. The main contributions of this work include:
@@
    We propose an LLM-in-the-loop clustering paradigm under both context-free and context-aware settings, which effectively utilizes intermediate results during the clustering process. This method surpasses state-of-the-art baselines and offers significant practical advantages in downstream applications.
@@ 2 Related Work
The discussion of text clustering with LLMs integration emerged as a trending topic since 2023. For instance, Zhang et al. (2023) utilized instruction-tuned LLMs like ChatGPT to better align clustering granularity with user preferences by refining the embedding space, while Viswanathan et al. (2024) demonstrated that LLMs can significantly enhance clustering quality by improving input features, providing constraints during clustering, and performing post-corrections. Feng et al. (2024) proposed a simple approach to refine edge points with LLM, resulting in minor performance improvement. Recently, Hong et al. (2025) proposed the concept of LLM-in-the-loop machine learning, categorizing LLM integration into data-, model-, and task-centric approaches. A “human-in-the-loop” solution can be replicated by LLM utilities Chen et al. (2024), enabling convenient and cost-efficient industrial deployment.
@@
Intent clustering is a fundamental step in intent induction Chandrakala et al. (2024) and the development of intent classifiers Gung et al. (2023). While various techniques have been proposed concerning semi-supervised Kumar et al. (2022) and deep learning approaches Lin et al. (2020); Zhang et al. (2021), recent advancements in LLM-based intent clustering systems have shown significant practical benefits Liang et al. (2024). Notably, the IDAS method proposed by De Raedt et al. (2023) is distinguished by its innovative use of LLMs and in-context learning to generate descriptive utterance labels, thereby enhancing clustering performance. However, the effective use of LLMs continues to pose challenges, such as managing input length and ensuring model robustness Song et al. (2023).
@@
Previous studies have primary focused on a data-centric approach, prioritizing the enhancement of input text representations. In contrast, this paper emphasizes on the importance of intermediate results during the clustering process, designing tailored solutions for human-aligned intent clustering. Additionally, rather than relying on public benchmark datasets with limited complexity, this work tackles a more realistic setting by clustering a noisy dataset transcribed from real customer service calls, offering practical solutions in handling dialogue intent clustering.
3 Preliminaries
3.1 Coherence Evaluation
Inspired by human behaviors in perceiving texts and making comparisons based on semantic meaning rather than word-level similarity Peter W. Foltz and Landauer (1998), this paper proposes the semantic consistency within a cluster, known as semantic coherence, as a superior metric and optimization target. The evaluation is simplified by formulating it as a binary classification problem, leading toward intuitive evaluation and easing the training of the LLM evaluator. Good clusters consist of sentences focused on a specific topic, while Bad clusters contain inconsistent or ambiguous intentions (see Table 6 for examples). This task is particularly challenging for traditional machine learning classifiers due to a lack of semantic understanding Mimno et al. (2011), necessitating fine-tuned LLMs Gu et al. (2025) for robust evaluation.
3.2 Cluster Naming
Giving each cluster a concise and meaningful name is essential for many downstream applications Pattnaik et al. (2024). In this paper, we introduce a novel naming convention, "Action-Objective," which is particularly effective for capturing dialogue intents that are typically topic-oriented (e.g., insurance, loan) and involve distinct actions (e.g., inquire). Examples of human annotations are shown in Table 7, and a comparison of different naming conventions is presented in Table 8.
@@ 4 Proposed Methods
4.1 Context-Free Intent Clustering
Context-free intent clustering represents the simplest setting, which ignores task-related information such as dialogue roles and domain-specific knowledge. The performance of the context-free approach provides a general expectation of effectiveness when applying the proposed methods to different datasets and tasks.
4.1.1 LLM-in-the-loop Intent Clustering
With the introduction of cluster-level semantic coherence evaluation, individual clusters that are identified to be badly clustered can be further refined, enabling the iterative discovery of high-quality clusters. The proposed method is presented in Algorithm 1. To determine the appropriate number of clusters, which is often unknown in real-world applications, a candidate space of cluster numbers $\bm{N}$ is predefined. In each iteration, text embeddings $\bm{t}$ are input into a clustering algorithm $\bm{F}(\bm{t},n_{i})=\bm{c_{n_{i}}}$ for each candidate cluster number $n_{i}$. The coherence of the clustering results is then evaluated by a fine-tuned LLM, and the cluster assignment yielding the highest number of good clusters is selected as the intermediate result. The good clusters in this result are retained, and the sentences in the bad clusters are re-clustered in the subsequent iteration, ensuring that the final result consists mostly of high-quality clusters. The final clusters are labeled with LLM naming to improve the interpretability and practical values.
Determining the optimal number of clusters in a noisy text corpus is often challenging. Here, we propose a local search heuristic that maximizes the “good/bad” ratio at each iteration:
$\max_{n_{k}\in\bm{N}}\quad\frac{\#\,\bm{c_{n_{k}}}|\,g_{k}=good}{\#\,\bm{c_{n_{k}}}|g_{k}=bad}$
where $n_{k}$ denotes the best cluster number at the $k$th iteration. This enables the automatic discovery of suitable cluster numbers in a step-by-step manner. The search space $\bm{N}$ should chosen carefully to balance accuracy and efficiency; thus, we propose a search space pruning strategy, as illustrated in Section 6.4.2 for more details.
@@ 4.1.1 LLM-in-the-loop Intent Clustering
1:sentences, $LLM_{eval}$, $LLM_{name}$, cluster numbers $\bm{N}=[n_{1},\dots,n_{k}]$
2:set of clusters $\bm{C}$ with intent labels $\bm{L}$
3:$\bm{t}\leftarrow\text{SentenceTransformer}(sentences)$
4:$\bm{C}\leftarrow\emptyset$
5:for each $n_{i}\in N$ do
6:     initial clusters $\bm{c_{n_{i}}}\leftarrow\bm{F}(\bm{t},n_{i})$
7:     $\bm{g_{i}}\leftarrow[LLM_{eval}(c_{1}),\dots,LLM_{eval}(c_{n_{i}})]$
8:end for
9:$i\leftarrow\mathop{\arg\max}_{i}\frac{\#\,\bm{c_{n_{i}}}|\,g_{i}=good}{\#\,\bm{c_{n_{i}}}|g_{i}=bad}$
10:$\bm{C}\leftarrow\bm{C}\oplus\bm{c_{n_{i}}}\;|\,g_{i}=good$
11:if #$\#\,\bm{c_{n_{i}}}|\,g_{i}=bad>\epsilon$ then
12:     reverse to 3 with $\bm{t}\leftarrow\bm{c_{n_{i}}}\;|\,g_{i}=bad$
13:end if
[…]
4.1.2 Post-Correction with Intent Labels
From an initial exploration, multiple clusters representing similar intents may emerge in the final result (see Table 9). This occurs because, as the iterative process progresses, the number of poorly clustered sentences decreases, resulting in smaller and less representative clusters being formed. The challenge of embedding distance prevents the grouping of these clusters, motivating post-correction by merging similar clusters based on intent labels.
Unlike Viswanathan et al. (2024) and Feng et al. (2024), which directly query LLMs the correctness of cluster assignments (i.e., whether the cluster assignment of a data point is correct), we leverage the cluster naming capability of LLMs to label each cluster with its underlying intent. These labels $L(c_{i})$ can be merged when $\text{Dist}(L(c_{i}),L(c_{j}))<\theta$, where $\theta$ is a predefined merge threshold and Dist is a distance metric. As the intent label is a concrete and concise summary of the whole cluster, the embeddings of these labels are believed to retain semantic information and are suitable for merging similar clusters.
4.2 Context-Aware Intent Clustering
From a practical perspective, dialogue datasets are often accompanied by domain-specific features or use-case scenarios (e.g., customer service calls, group discussions). This motivates the incorporation of contextual information into the clustering process. However, the discovery of suitable knowledge remains a challenging task; thus, a knowledge discovery process is proposed based on LLM naming, where intermediate clustering assignments are utilized to achieve improved clustering results.

### v3

@@ Abstract
Discovering customer intentions in dialogue conversations is crucial for automated service agents. However, existing intent clustering methods often fail to align with human perceptions due to a heavy reliance on embedding distance metrics and a tendency to overlook underlying semantic structures. This paper proposes an LLM-in-the-loop (LLM-ITL) intent clustering framework, integrating the semantic understanding capabilities of LLMs into conventional clustering algorithms. Specifically, this paper (1) investigates the effectiveness of fine-tuned LLMs in semantic coherence evaluation and intent cluster naming, achieving over 95% accuracy aligned with human judgments; (2) designs an LLM-ITL framework that facilitates the iterative discovery of coherent intent clusters and the optimal number of clusters; and (3) proposes context-aware techniques tailored for customer service dialogue. As existing English benchmarks offer limited semantic diversity and intent groups, we introduce a comprehensive Chinese dialogue intent dataset, comprising over 100k real customer service calls and 1,507 human-annotated intent clusters. The proposed approaches significantly outperform LLM-guided baselines, achieving notable enhancements in clustering quality and lower computational cost. Combined with several best practices, our findings highlight the potential of LLM-in-the-loop techniques for scalable and human-aligned intent clustering.
@@
Mengze Hong1, Wailing Ng1, Chen Jason Zhang1 Yuanfeng Song2, Di Jiang2thanks: Corresponding author: dijiang@webank.com 1Hong Kong Polytechnic University 2AI Group, WeBank Co., Ltd
@@ 1 Introduction
Figure 1: Comparison of embedding-distance metric and human perception: cosine similarity failed to identify the same intentions under diverse expression (top), and to distinguish distinct intentions under similar expressions (bottom).
Intent discovery is a crucial task in numerous NLP applications, including the design of AI-driven dialogue systems Hengst et al. (2024) and the learning of utterance patterns Ghosal et al. (2020). While intent clustering techniques are extensively studied to automatically identify thematic relationships within text corpora Gung et al. (2023), existing research primarily focuses on developing meaningful sentence representations and relies on embedding distance metrics for optimization Yin et al. (2021). This approach often overlooks the distinctive characteristics of textual information, such as linguistic patterns and semantic diversity (see Figure 1), and restricts human-aligned evaluation and validation of clustering quality Vinh et al. (2009). This issue is particularly pronounced in languages with rich semantics, such as Chinese, where two seemingly similar sentences can convey entirely different meanings, and vice versa Chen (1993).
With growing research interest in integrating Large Language Models (LLMs) into the problem-solving pipeline Hong et al. (2025), LLM-guided clustering techniques have emerged, demonstrating superior performance over traditional machine learning algorithms. While these methods effectively incorporate language understanding into the clustering process, they primarily focus on data preprocessing, querying LLMs for embedding refinement or data augmentation Zhang et al. (2023); Viswanathan et al. (2024). This approach represents a surface-level integration, potentially missing the benefits of LLMs in providing semantic-driven guidance to the clustering process and the quality of the final results.
In this paper, we introduce an LLM-in-the-loop (LLM-ITL) intent clustering method, designed to facilitate the iterative discovery of coherent intent clusters. Our approach adeptly utilizes intermediate clustering results by incorporating human-aligned LLM utilities, offering computationally efficient and human-interpretable intent clustering. The key contributions of this work are:
@@
    We propose an LLM-in-the-loop clustering algorithm to derive an optimal set of clusters in scenarios where the number of clusters is unknown. This approach effectively leverages intermediate results generated during the clustering process, outperforming state-of-the-art baselines and providing practical benefits for downstream applications. Additionally, the discussion on data sampling and LLM-based crowdsourcing further improves performance.
@@ 2 Related Work
The integration of LLMs into text clustering has become increasingly prominent since 2023. Zhang et al. (2023) introduced ClusterLLM, an innovative approach that leverages instruction-tuned LLMs like ChatGPT to refine sentence embedding spaces through pairwise preference questions, aligning clustering granularity with user preferences. Viswanathan et al. (2024) highlighted the enhancement of clustering quality by LLMs through feature improvement, the imposition of constraints during clustering, and post-correction processes. Additionally, Feng et al. (2024) proposed refining edge points with LLMs, which led to notable performance gains. Recently, Hong et al. (2025) conceptualized LLM-in-the-loop machine learning, classifying LLM integration into data-, model-, and task-centric strategies. This method mirrors a “human-in-the-loop” approach Chen et al. (2024), facilitating effective LLM integration in machine learning applications.
@@
Intent clustering is a fundamental step in intent induction Chandrakala et al. (2024) and the development of intent classifiers Gung et al. (2023). While semi-supervised Kumar et al. (2022) and deep learning methods Lin et al. (2020); Zhang et al. (2021) have significantly advanced intent clustering, recent progress in LLM-based systems has highlighted their potential practical advantages Liang et al. (2024). The IDAS method effectively employs in-context learning to generate descriptive utterance labels and enhance sentence embeddings De Raedt et al. (2023). However, challenges such as computational cost and robustness remain unresolved Song et al. (2023).
@@
Previous studies have primarily focused on enhancing input text representations and the direct use of LLMs for cluster modification, imposing high computational cost and uncertainty. In contrast, this paper implements human-aligned LLM utilities to create a LLM-in-the-loop intent clustering process with emphasis on the intermediate results. Additionally, this work tackles a more realistic scenario by releasing a complex dataset transcribed from real customer service calls, offering practical insights and contributing to more meaningful evaluations in future intent clustering.
Figure 2: Overview of proposed LLM-in-the-loop Dialogue Intent Clustering.
@@ 3 Proposed Methods
In this section, we first outline the design of human-aligned LLM utilities. Then, we introduce a comprehensive LLM-in-the-loop intent clustering framework (see Figure 2) and show how each LLM utility contributes to the clustering process.
3.1 LLM Utilities
Coherence Evaluation.
Inspired by human behaviors in perceiving texts and making comparisons based on semantic meaning rather than word-level similarity Peter W. Foltz and Landauer (1998), this paper proposes the semantic consistency within a cluster, known as semantic coherence, as a superior metric and optimization target. The coherence evaluation is formulated as a binary classification problem, leading toward intuitive evaluation and easing the training of the LLM evaluator (see Section 5.6 for different formulations). Good clusters consist of sentences focused on a specific topic, while Bad clusters contain inconsistent or ambiguous intentions (see Table 9 for examples). This task is particularly challenging for traditional machine learning classifiers due to a lack of semantic understanding Mimno et al. (2011), necessitating fine-tuned LLMs Gu et al. (2025) for robust evaluation (see Section 5.1 for fine-tuning details).
Cluster Naming.
Giving each cluster a concise and meaningful name is essential for many downstream applications Pattnaik et al. (2024). In this paper, we introduce a novel naming convention, "Action-Objective," which is particularly effective for capturing dialogue intents that are typically topic-oriented (e.g., insurance, loan) and involve distinct actions (e.g., inquire, confirm). Examples of human annotations are shown in Table 10, and a comparison of different naming conventions is presented in Table 11, demonstrating the superiority of the proposed convention in capturing customer intention with clear goal-oriented expression.
3.2 LLM-in-the-loop Iterative Intent Clustering with Coherence Evaluation
@@
1:Unlabeled sentence corpus $\mathcal{S}$, embedding function $f_{\text{emb}}$, coherence evaluator $\mathcal{M}_{\text{eval}}$, candidate cluster numbers $N=\{n_{1},\dots,n_{k}\}$, threshold $\epsilon>0$, max iterations $T_{\text{max}}$
2:Set of clusters $\mathcal{C}$
3:$\mathcal{E}\leftarrow\{f_{\text{emb}}(s)\mid s\in\mathcal{S}\}$
4:$\mathcal{S}^{(0)}\leftarrow\mathcal{S}$, $\mathcal{C}\leftarrow\emptyset$, $t\leftarrow 0$
5:while $\frac{|\mathcal{S}^{(t)}|}{|\mathcal{S}|}>\epsilon$ and $t<T_{\text{max}}$ do
6:     $\mathcal{E}^{(t)}\leftarrow\{f_{\text{emb}}(s)\mid s\in\mathcal{S}^{(t)}\}$
7:     for each $n_{i}\in N$ do
8:         $\mathcal{C}_{n_{i}}^{(t)}\leftarrow F(\mathcal{E}^{(t)},n_{i})$
9:         $\mathbf{g}_{n_{i}}^{(t)}\leftarrow[\mathcal{M}_{\text{eval}}(C_{1}^{(t)}),\dots,\mathcal{M}_{\text{eval}}(C_{n_{i}}^{(t)})]$
10:     end for
11:     $n^{*}\leftarrow\arg\max_{n_{i}\in N}\frac{\sum_{j=1}^{n_{i}}\mathbb{I}\left(\mathbf{g}_{n_{i}}^{(t)}[j]=1\right)}{\sum_{j=1}^{n_{i}}\mathbb{I}\left(\mathbf{g}_{n_{i}}^{(t)}[j]=0\right)+1}$
12:     $\mathcal{C}_{\text{good}}^{(t)}\leftarrow\{C_{j}\in\mathcal{C}_{n^{*}}^{(t)}\mid\mathbf{g}_{n^{*}}^{(t)}[j]=1\}$
13:     $\mathcal{C}\leftarrow\mathcal{C}\cup\mathcal{C}_{\text{good}}^{(t)}$
14:     $\mathcal{S}^{(t+1)}\leftarrow\mathcal{S}^{(t)}\setminus\bigcup_{C\in\mathcal{C}_{\text{good}}^{(t)}}C$
15:     $t\leftarrow t+1$
16:end while
17:return $(\mathcal{C})$
At iteration $t$, the current set of unassigned sentences is denoted as $\mathcal{S}^{(t)}$. A special case in the first iteration, where $\mathcal{S}^{(0)}=\mathcal{S}$ represents the entire set of unique sentences derived from the dialogue corpus. For each candidate cluster number $n_{i}\in N$, we compute:
$\mathcal{C}_{n_{i}}^{(t)}=F(\mathcal{E}^{(t)},n_{i}),$
where $F$ is a clustering function (e.g., K-means clustering), and $\mathcal{E}^{(t)}=\{\mathbf{e}_{s}\mid s\in\mathcal{S}^{(t)}\}$ representing the sentence embeddings. This results in $N$ distinct cluster assignments.
Then, the semantic coherence of each cluster is evaluated using a fine-tuned LLM $\mathcal{M}_{\text{eval}}$:
$\mathbf{g}_{n_{i}}^{(t)}=\left[\mathcal{M}_{\text{eval}}(C_{1}^{(t)}),\ldots,\mathcal{M}_{\text{eval}}(C_{n_{i}}^{(t)})\right],$
where $\mathcal{M}_{\text{eval}}(C_{k})=1$ if the cluster is coherent (i.e., “good” cluster), else 0.
While the number of clusters is often known with benchmark evaluation, it is challenging but necessary to implement mechanisms for determining the optimal number of clusters in a noisy text corpus. Here, we propose a local search heuristic that maximizes the “good/bad” ratio at each iteration. The optimal $n_{t}^{*}$ is selected as:
$n_{t}^{*}=\arg\max_{n_{i}\in N}\frac{\sum_{j=1}^{n_{i}}\mathbb{I}\left(\mathbf{g}_{n_{i}}^{(t)}[j]=1\right)}{\sum_{j=1}^{n_{i}}\mathbb{I}\left(\mathbf{g}_{n_{i}}^{(t)}[j]=0\right)+1},$
where $n_{t}$ denotes the best cluster number at iteration $t$. This enables the automatic discovery of suitable cluster numbers in a step-by-step manner. The search space $\bm{N}$ should be chosen carefully to balance accuracy and efficiency; thus, we propose a search space pruning strategy, as illustrated in Section 5.6 for more details.
Finally, the “good” clusters in the optimal $\mathcal{C}_{n^{*}}^{(t)}$ are retained, and the remaining sentences will be refined in the next (i.e., $t+1$) iteration, enabling the iterative discovery of high-quality clusters. The proposed method is summarized in Algorithm 1, and the LLM integration is highlighted.
3.3 Post-Correction with LLM-Generated Intent Labels
Preliminary results in Table 12 suggest that multiple clusters may capture similar intents due to diminishing size of unassigned sentences $\mathcal{S}^{(t)}$ in later iterations, leading to smaller and less representative clusters. The embedding distances between sentences within clusters limit the natural consolidation of clusters with similar intents but different expressions Khan et al. (2020), necessitating a post-correction step to merge semantically aligned clusters. In contrast to previous methods that rely on direct LLM queries for validating individual cluster assignments Viswanathan et al. (2024); Feng et al. (2024), which are computationally expensive and inconsistent, we propose a context-aware approach, leveraging LLMs’ naming utility to robustly merge clusters based on intent labels.
At the end of the iterative clustering process, each cluster $C_{k}$ receives an intent label from the fine-tuned LLM naming utility:
$l_{k}=\mathcal{M}_{\text{name}}(C_{k}).$
These labels concisely summarize the semantic information of each cluster and are mapped to the semantic space of sentence embeddings using the embedding function $f_{\text{emb}}$:
$\mathbf{l}_{k}=f_{\text{emb}}(l_{k})\in\mathbb{R}^{d},\quad\|\mathbf{l}_{k}\|_{2}=1.$
This normalization positions label embeddings $\mathbf{l}_{k}$ on the unit hypersphere $\mathbb{S}^{d-1}=\{\mathbf{x}\in\mathbb{R}^{d}\mid\|\mathbf{x}\|_{2}=1\}$, enabling accurate measurement of semantic relationships along the sphere’s surface rather than through straight-line distances Fletcher et al. (2004), thus facilitating precise intent similarity comparisons in high-dimensional space.
Then, the clusters are structured into a semantic affinity graph $G=(V,E)$ to model relationships based on label similarity. Vertices $V=\{C_{1},C_{2},\ldots,C_{K}\}$ represent clusters, and the edges $E\subseteq V\times V$ are determined using hyperspherical geometry by computing the geodesic distance, which captures angular separation between label embeddings Fletcher et al. (2004):
$\text{Dist}(\mathbf{l}_{i},\mathbf{l}_{j})=\arccos(\langle\mathbf{l}_{i},\mathbf{l}_{j}\rangle),$
forming edges if $\text{Dist}(\mathbf{l}_{i},\mathbf{l}_{j})<\theta$, with $\theta$ is a predefined threshold ($\theta=0.8$).
For each derived edge, a probabilistic criterion is designed to enhance the robustness of merging decisions, naturally modeling label embeddings as samples from a mixture of von Mises-Fisher distributions Banerjee et al. (2005):
$p(\mathbf{l}_{k}\mid\boldsymbol{\mu}_{m},\kappa_{m})=Z_{d}(\kappa_{m})\exp(\kappa_{m}\langle\mathbf{l}_{k},\boldsymbol{\mu}_{m}\rangle),$
where $\boldsymbol{\mu}_{m}=\mathbf{l}_{m}\in\mathbb{S}^{d-1}$ is the mean direction of the $m$-th intent embedding, $\kappa_{m}>0$ controls the distribution’s tightness, and the normalization constant $Z_{d}(\kappa_{m})$ ensures the density integrates to 1 over the hypersphere:
$Z_{d}(\kappa)=\frac{\kappa^{d/2-1}}{(2\pi)^{d/2}I_{d/2-1}(\kappa)}.$
Thus, the probability that clusters $C_{i}$ and $C_{j}$ share the same intent is computed for each edge as:
$P(\text{same}\mid\mathbf{l}_{i},\mathbf{l}_{j})=\sum_{m=1}^{K}\pi_{m}p(\mathbf{l}_{i}\mid\boldsymbol{\mu}_{m},\kappa_{m})p(\mathbf{l}_{j}\mid\boldsymbol{\mu}_{m},\kappa_{m}),$
with $\pi_{m}=1/K$ as uniform mixture weights. An edge is retained only if this probability exceeds a predefined threshold $\tau$ ($\tau=0.7$), minimizing inappropriate merges by accounting for uncertainty.
Finally, clusters linked in the affinity graph are consolidated into connected components, forming a refined cluster assignment $\mathcal{C}^{\prime}=\{C^{\prime}_{1},C^{\prime}_{2},\ldots,C^{\prime}_{K^{\prime}}\}$ that eliminates redundancy, enhances interpretability, and maintains robust semantic alignment. Each merged cluster $C^{\prime}_{k}$ receives a new intent label reflecting its semantic content.
3.4 Context-Aware Role Separation with LLM-Generated Intent Labels